OpenStudy (anonymous):

The following geometric sequences represent the populations of two bacterial cultures at the 1-hour mark, the 2-hour mark, the 3-hour mark, and so on. Culture A starts with more bacteria, but culture B has a ratio of increase that is larger. Which culture will have the greatest population at the 19-hour mark? A. 800, 1,200, 1,800, 2,700, ... B. 5, 10, 20, 40, ...

OpenStudy (ddcamp):

Every hour, A is increased by 1.5 times, and B is doubled. Using this we can say that A(t) = 800 * (1.5)^(t-1) B(t) = 5 * (1.5)^(t-1) Where A(t) is the size of colony A at time t and B(t) is the size of colony B at time t.

OpenStudy (anonymous):

so my answer is A! thanks!